An airplane travels at 450 km/he. it can fly 20.0 miles on one gallon of fuel. the density of the fuel is 0.782g/mL. How many pounds of fuel will this plane need to travel from Chicago to Miami, a distance of 1400 miles?
1
Expert's answer
2013-09-10T08:29:59-0400
1 pound = 0.453 kg
1 gallon = 3.785 L
0.782 g/mL = 0.782 kg/L
So an airplane for 20.0 miles uses 1 gallon = 3.785 L of fuel.
For a distance from Chicago to Miami (1400 miles) it needs: 1400/20*3.785 = 264.95 L of fuel.
According to the indicator of density, the mass will be: 264.95*0.782 = 207.2 kg
According to the previous data (1 pound = 0.453 kg) the needed mass in pounds will be: 207.2/0.453 = 457.4 pounds